# Human-AI Collaboration: Trade-offs Between Performance and Preferences

### v1

@@ 2.4.2 Trade-off Between Performance and Preference
3 Discussion
@@
    People prefer AI collaborators that enable meaningful human contribution. In our study, collaborative agents that showed greater performance differences compared to the participants were generally less preferred. This observation aligns with previous research, which demonstrated that inequity aversion can enable groups of autonomous agents to maintain cooperative behavior [23]. However, our study is the first to demonstrate this effect in a human-AI collaboration context. We also found that people are more averse to falling behind the AI than they are to getting ahead of the AI. This tendency highlights some asymmetry that could reflect a human preference to lead the human-AI team.
@@
    People prefer AI collaborators that are considerate of human intentions. Participants showed a clear aversion to perceived intrusions into the tasks they delegated for themselves. Likewise, participants’ preferences were not significantly influenced by the AI’s actual performance or their subjective assessment of its performance. Our findings also demonstrate how a considerate, or human-aware, AI collaborator does not always take away from the objective performance of the human-AI team. We even found that considerate AI collaborators contribute to collective performance as well as egocentric AI collaborators. For example, our results showed that the human-Omit team performed as well as, and sometimes better than, the team that paired a human with a performance-maximizing agent. This early evidence of complementarity in human-AI collaboration—where human-aware algorithms boost individual and team performance—suggests that more sophisticated human-aware AI systems may exceed the performance of teams paired with purely performance-optimized agents. Taken together, our findings motivate reasons to design collaborative AI that better supports human capabilities, as this may increase user adoption and human-AI team performance.
@@
3.1 Limitations and Future Work
@@ 4 Conclusion
According to our analysis, people prefer to collaborate with algorithms that enable the user to contribute meaningfully to the team, with an AI teammate that complements their actions rather than dominating the interaction. We found evidence of preferences being informed by a tendency to favor inequity aversion, implying that the collaborative AI’s capabilities should be comparable to those of the human user. We also found strong support for preferences in agents that show a tendency to defer to the human. One avenue for achieving such approximations of collaborative abilities is through labor divisions that emphasize user autonomy, for example, through task delegation or spatial separation. Our experimental paradigm can be leveraged to iteratively improve the collaborative abilities of new and existing systems alike.
5 Methods
5.1 Participants
5.2 Game Environment
There are two agents in the game: a collaborative AI player and a human player, each of whom have their own unique icon and distinctively colored square. The objective of the game is to intercept as many moving, point-valued, targets as possible within a fixed time frame. Both the human player and the AI player can intercept targets but each target can be intercepted only once. Optimal task performance necessitates quick strategic decisions from the human player to intercept targets in a particular sequence during the limited time they are available while also paying attention to and coordinating with the actions of the collaborative AI player.
Targets spawn randomly at the edge of the circular game area. Their initial movement angle is randomly set within a cone of possible angles. Spawned targets move in straight-line directions at constant speed, sampled from a uniform distribution that ensures target speeds are between 1-50% slower than the player’s avatar. Targets exit the playable area if they are not intercepted. The spawning process ensures that the number of objects in the game area is limited to the target density (either 5 or 15 targets). One key feature of the spawning process is that it is independent of the player’s skill in interception. After a player intercepts a target, it disappears from the game area, but its path is still computed until it hits the perimeter. Only once a target, visible or not, hits this perimeter will a new target be spawned. Each target is worth between zero and fifteen points, with the probability distribution of point values following a Beta(1,2) distribution that we discretized over point-values via binning. In practice, this means that low-value targets appear more often than high-value targets.
5.3 Collaborative AI Agents
5.3.1 Search Algorithm
The search algorithm is designed to approximate optimal solutions to the target interception task. The algorithm computes all possible interception sequences involving up to three targets, updating the positions of both the AI player and the targets throughout the sequence. This ensures that the AI player can respond to dynamic changes in the game state. The three-target limit is imposed to ensure the algorithm can operate in real time during behavioral experiments. For more details, see Appendix 7.2.
5.3.2 Agent Variations
The search algorithm formed the basis for all agents in our experiment. We developed several variations of the search algorithm to create different collaborative AI players. These variations included changes to the target consideration set (which targets the AI player could pursue), delays in initiating a plan, and perception of point values. Variations were conceived as mechanisms that incorporate and give rise to heightened perceptions of traits observed in previous research [5, 10]. This study is not designed to test all possible combinations of these features. Instead, we focused on a set of five agents that test out a key set of variations:
  1. 1.
  2. 2.
  3. 3.
    Divide Agent: This agent operates in the same fashion as the omit agent but applies a divide and conquer strategy. This was done to make it easier for the human player and AI agent to avoid getting in each other’s way and, potentially, have better task delegation [25, 7]. The agent (virtually) divides the game area in two halves where the dividing line is orthogonal to the imaginary line from the human player to the game’s center-point. The agent only considers targets that can be intercepted in the half not occupied by the human player, with the allotted area being continuously recomputed as the human’s position changes. Therefore, this agent omits a larger set of targets from consideration than the omit agent, leaving more targets for the human player, further reducing potential interception conflicts between the players.
  4. 4.
  5. 5.
5.4 Procedure
Figure 9: Illustration of the procedure. Top and bottom rows (1 and 2) illustrate the two blocks in the experiment. Within each block, participants play two rounds, one with each AI agent from their assigned pair. Agents are then evaluated on a variety of dimensions using 7-point Likert scales. After submitting their ratings, participants indicate their preferred agent in a two-alternative forced choice. Finally, participants are asked to provide free-text responses explaining why they chose the agent they preferred. This procedure is repeated over two blocks where target density is varied. In the illustration, the first and second block have low and high target density. The density order is counterbalanced in the experiment.
5.5 Design
5.6 Data Analysis
To assess statistical significance, we utilized Bayes factors (BFs) to determine the extent to which the observed data adjust the a priori belief in the alternative and null hypotheses. Values of 3 $<$ BF $<$ 10 and BF $>$ 10 indicate moderate and strong evidence against the null hypothesis, respectively. Similarly, values of 1/10 $<$ BF $<$ 1/3 and BF $<$ 1/10 indicate moderate and strong evidence in favor of the null hypothesis, respectively [27, 28, 29].
Estimation of the logistic regression in Equation 2 was performed with Bayesian methods in the JASP (Version 0.19, [31]) environment using the default priors based on the Generalized g-Prior Distribution (CCH; [32]) with $\alpha=0.5$, $\beta=2$, and $s=0$. In addition to Bayes factors for each individual covariate, we also report the 95% credible interval (CI). Although it might be tempting to use the CI to test hypotheses (e.g., rejecting the null hypothesis if the CI does not include the null value), in accordance with recent recommendations [33, 34] we use a more conservative approach, where the CI becomes relevant only after the BF shows evidence for the alternative hypothesis.
6 Declarations
@@
NA
Data and materials availability
The behavioral datasets generated during the current study are available at OSF, https://osf.io/ybweq/?view_only=cb4d4c7ac0b848b79b6ae8c7b09278cc
Code availability
@@
LM, SK, JA, MS, YY, DT, and MS designed the study. SK developed and ran the behavioral experiment. LM, SK, JA, and MS analyzed the data. LM, SK, and MS were major contributors in writing the manuscript. All authors read and approved the final manuscript.

### v2

@@ Results
Figure 10: Scatter plot of team performance over probability of being chosen, by AI agent condition and target density conditions. Error bars reflect the standard error from the mean. Results are from Experiment 1.
Experiment 2
The results of Experiment 1 indicated that participants’ choices of collaborative AI agents were influenced more strongly by subjective impressions than by either their own performance or the AI’s performance. One possibility is that certain aspects of the experimental design contributed to this outcome. In Experiment 1, after participants interacted with two agents, they were first presented with a questionnaire assessing the agents on multiple collaborative dimensions, and only afterward were they asked to choose their preferred agent. This ordering may have shaped participants’ subsequent choices by drawing attention to the specific traits emphasized in the questionnaire, thereby amplifying the influence of subjective factors relative to objective performance metrics. Additionally, because monetary compensation was fixed and independent of performance, participants may have been less motivated to prioritize performance-related considerations when selecting an agent.
Experiment 2 was designed to address these potential limitations by introducing two changes. First, we manipulated the order in which participants completed the preference choice and the questionnaire. This allowed us to test whether the act of rating an agent on specific traits before making a choice alters the relative weight of subjective and performance-based factors. Second, we introduced a performance-based incentive: participants received a \$2 bonus if their team’s cumulative score across the four game rounds ranked in the top 50% of all participants. This bonus was intended to increase the salience of performance outcomes and test whether financial incentives would shift preferences toward higher-performing agents.
Methods
The methodological framework for Experiment 2 closely mirrored that of Experiment 1, with two key modifications: (1) the introduction of a within-subject manipulation of the choice–questionnaire order, and (2) the addition of a performance-based bonus incentive.
Participants
93 participants were recruited from the online participant recruitment platform Prolific. Ages ranged from 23 to 72 years (Mean = 43.36, SD = 11.72), with 42% identifying as female, 57% as male, and 1% choosing not to disclose gender. All participants resided in the United States and had not participated in Experiment 1. The study was conducted remotely on participants’ personal computers. Each participant received a base payment of 5 USD for completing the approximately 25-minute study, corresponding to an average base rate of 12.43 USD/hour. In addition, participants were eligible for a 2 USD bonus if their cumulative team score across the four main rounds of gameplay ranked in the top 50% of all participants within their assigned conditions. In total, 40 participants received bonus payments, bringing the final average hourly compensation rate to 14.38 USD/hour.
Procedure and Design
The game environment and agents were identical to Experiment 1. The study followed the same experimental manipulations and procedures as in Experiment 1.
The key change was the order manipulation, implemented within participants across the two blocks: each participant experienced both the Survey→Choice (after one block) and the Choice→Survey order (in the other block). Participants were randomly assigned to one of two conditions. In condition 1, participants completed Survey→Choice after Block 1 and Choice→Survey after Block 2 (N = 48), whereas participants in condition 2 completed Choice→Survey after Block 1 and Survey→Choice after Block 2 (N = 45). Thus, order varied within participants (by block), and the between-participant factor was which order occurred first. All other procedures—including instructions, tutorial and comprehension checks, rating scales, free-text rationales, and end-of-study feedback—were unchanged from Experiment 1.
Results
The analyses of Experiment 2 examined whether altering the order of the choice and questionnaire tasks influenced the relative weight of performance-based (objective) versus perception-based (subjective) factors in participants’ preferences, and whether the introduction of a performance-based bonus shifted these preferences. For additional experimental results, see Appendix F.
Predictive Models for Human Preferences
We applied the same Bayesian logistic regression framework used in Experiment 1 to model choice behavior. Models were fit separately for each order—Survey→Choice and Choice→Survey—with objective and subjective predictors analyzed independently.
As shown in Figure 11, there were no meaningful differences between the two order conditions in the predictive importance of either objective or subjective measures. In other words, the ordering of the survey and forced-choice tasks did not substantially change the relative influence of performance-based or subjective factors on agent preference.
One notable difference between the conditions emerged in the bias term. When the survey preceded the choice, participants were more likely to select the first agent presented (consistent with Experiment 1). When the choice preceded the survey, however, participants tended to select the second (most recently presented) agent. This pattern is consistent with a primacy/recency effect [18]: if the choice immediately follows the second agent’s presentation, recency may drive preference, whereas an intervening activity—such as answering survey questions—appears to shift attention back toward the first-presented agent.
Figure 11: Posterior ($\beta$) coefficients of Bayesian logistic regression models predicting choice. The coefficients are shown across objective and subjective metrics (left and right panels) and survey order (indicated by colors). Coefficient estimates can be thought of as weights for the importance of metrics in explaining choices. Error bars represent 95% credible intervals.
Predictive Accuracy
For the survey→choice condition, the model achieved predictive accuracies of 62% for objective metrics and 78% for subjective metrics—closely matching the performance levels observed in Experiment 1 (62% and 84%, respectively). When the order was reversed (choice→survey), predictive accuracies were nearly identical: 62% for objective metrics and 79% for subjective metrics.
Preferences for Collaborative Agents
To evaluate whether the introduction of performance-based incentives in Experiment 2 shifted participants’ preferences for collaborative agents, we compared marginal choice percentages with those observed in Experiment 1. In Experiment 2, agents were chosen in the following order from most to least preferred: Omit (74%), Delay (43%), Divide (42%), Bottom-Feeder (59%), and Ignorant (31%). In Experiment 1, the order was Omit (64%), Divide (59%), Delay (58%), Bottom-Feeder (49%), and Ignorant (22%). The overall ranking of agents was similar across experiments, with the only change being a swap in relative positions of the Delay and Divide agents, whose marginal preference rates were closely matched in both studies.
Discussion
The findings from Experiment 2 indicate that the ordering of choice and survey responses had negligible influence on the predictive power of either objective or subjective measures. The stability of these accuracies across conditions suggests that participants’ evaluation strategies and choice behaviors were not meaningfully altered by whether they rated the agents before or after making their selection. In both orderings, subjective ratings remained substantially more predictive of choice than objective performance metrics, consistent with the patterns observed in Experiment 1.
In addition, the absence of an incentive effect in this study suggests that, under the modest stakes offered here, participants’ choices of collaborative agents were not influenced by the prospect of a performance-based bonus. Even when a financial reward was contingent on achieving higher team scores, participants continued to prioritize agents they perceived as supportive, complementary, and considerate of their actions over those that simply maximized team performance. However, it remains possible that larger or more salient incentives could shift these priorities, leading participants to place greater weight on performance outcomes when selecting collaborators. Future work should focus on high-stakes settings to determine whether and how stronger incentives alter human-AI preference trade-offs.
General Discussion
@@ Discussion
    People prefer AI collaborators that enable meaningful human contribution. In our study, collaborative agents that showed greater performance differences compared to participants were generally less preferred. This observation aligns with previous research, which demonstrated that inequity aversion can enable groups of autonomous agents to maintain cooperative behavior [13]. However, our study is the first to demonstrate this effect in a human-AI collaboration context. We also found that people are more averse to falling behind the AI than they are to getting ahead of the AI. This tendency highlights some asymmetry that could reflect a human preference to lead the human-AI team.
@@
    People prefer AI collaborators that are considerate of human intentions. Participants showed a clear aversion to perceived intrusions into the tasks they delegated for themselves. Likewise, participants’ preferences were not significantly influenced by the AI’s actual performance or their subjective assessment of its performance. Our findings also demonstrate how a considerate, or human-aware, AI collaborator does not always take away from the objective performance of the human-AI team. We even found that considerate AI collaborators contribute to collective performance as well as egocentric AI collaborators. For example, our results showed that the human-Omit team performed as well as, and sometimes better than, the team that paired a human with a performance-maximizing agent. This evidence of complementarity in human-AI collaboration—where human-aware algorithms boost individual and team performance—suggests that more sophisticated human-aware AI systems may exceed the performance of teams paired with purely performance-optimized agents. Taken together, our findings motivate reasons to design collaborative AI that better supports human capabilities, as this may increase user adoption and human-AI team performance.
@@
  5. 5.
    People’s preferences for AI collaborators remain stable even when modest performance incentives are introduced. In Experiment 2, introducing a performance-based bonus and varying the order of preference and survey tasks did not meaningfully shift the balance between subjective impressions and performance metrics in predicting choice. Participants continued to favor agents perceived as supportive and considerate of their actions over those that maximized team scores. These findings suggest that, under low-stakes conditions, human-centered design features may outweigh performance considerations in shaping AI collaborator preferences.
Limitations and Future Work
@@ Conclusion
According to our analysis, people prefer to collaborate with algorithms that enable them to contribute meaningfully to the team, with an AI teammate that complements their actions rather than dominating the interaction. We found evidence of preferences being informed by a tendency to favor inequity aversion, implying that the collaborative AI’s capabilities should be comparable to those of the human user. We also found strong support for preferences in agents that show a tendency to defer to the human. One avenue for achieving such approximations of collaborative abilities is through labor divisions that emphasize user autonomy, for example, through task delegation or spatial separation. Our experimental paradigm can be leveraged to iteratively improve the collaborative abilities of new and existing systems alike.
Declarations
@@
Availability of data and materials
The behavioral datasets generated during the current study are available at OSF, https://osf.io/ybweq/?view_only=cb4d4c7ac0b848b79b6ae8c7b09278cc.
The experiment code is available on GitHub, https://github.com/MADLABatUCI/target-intercept-collab-ai.
Open Practices Statement
The behavioral datasets generated during the current study are available at OSF, https://osf.io/ybweq/?view_only=cb4d4c7ac0b848b79b6ae8c7b09278cc.
The experiments reported in this article were not pre-registered.
@@
Not applicable
@@
LM, SK, JA, MS, YY, DT, and MS designed the study. SK developed and ran the behavioral experiments. LM, SK, JA, and MS analyzed the data. LM, SK, and MS were major contributors in writing the manuscript. All authors read and approved the final manuscript.
Acknowledgments
Not applicable
Appendix A Questionnaire
Table 2: Survey Questions
Q1 “The bot and I were a team.”
Q2 “The bot was competent.”
Q3 “I understood the bot’s intentions.”
Q4 “The bot understood my intentions.”
Q5 “I contributed more to the team’s performance.”
Q6 “The bot was easy to play with”
Q7 “The bot was fun to play with.”
Q8 “The bot and I had a similar playing style.”
Appendix B Search Algorithm
The search algorithm computes all possible interception sequences involving up to three targets, updating the positions of both the AI player and the targets throughout the sequence. For each sequence, a total value is calculated by adding the intercepted objects’ points. The interception plan selected is based on the sequence with the highest cumulative point gain.
The planning algorithm incorporates a heuristic to account for the possibility of future objects entering the scene. More specifically, when planning for the first, second, and third interception, the algorithm discounts the value of future interceptions by a factor of $\alpha^{K}$ where $K$ is the estimated number of new objects that will enter the scene by the time a new interception is planned and $\alpha$ is a discounting parameter. Simulations determined that $\alpha=0.9$ led to good performance levels across different target densities. The search algorithm recomputes the plan in real time, allowing for changes in the game state, such as when a new target enters the game view or when a human player intercepts a target (preventing the AI player from intercepting it).
The algorithm includes a stability parameter to prevent the AI player from making erratic moves in response to new incoming targets. The stability causes the AI player to only change its current plan if a new plan is at least 20% better in expected point value than the current one. This threshold was determined through pilot studies, for which we found that lower thresholds resulted in participants perceiving the AI as overly erratic. Conversely, higher thresholds led the AI to appear overly rigid in its decision-making. By incorporating this stability component, the AI’s behavior becomes more predictable, allowing human players to plan around the AI player.
Appendix C Additional Results Experiment 1
Figure C1: Mean pixel distance between the human and AI player separated by AI agent type and target density. Gray shading visualizes the distribution of individual mean pixel distances, while error bars show the standard error of the mean.
Figure C2: Number of stealing occurrences for each player split by AI agent type and target density. A steal is defined as an instance where one agent marked a target before the other agent with the latter agent intercepting that target.
Appendix D Analysis of Open-Ended Responses Experiment 1
At the end of each experimental block, participants provided free-response feedback to identify aspects they liked or disliked about the AI agent’s performance. The goal of the analysis in this section is to understand if participants’ open-ended responses share characteristics that are present in the statements featured in the Likert questionnaire.
We first created single-word labels of participants’ descriptions using an approach that combined human raters and natural language processing (NLP). This analysis began with four raters classifying the comments as either positive or negative. Our four human raters then described the agents using a single word or short phrase. We then utilized an NLP procedure to ensure that words that appeared repeatedly were standardized to a single term, and multi-word aspects were condensed into concise, single-word labels. We then applied human validation to ensure that these single-word labels were properly balanced into a positive or negative classification.
Subsequently, three raters categorized these extracted terms according to the eight items on the evaluation scale, allowing us to compare participants’ subjective feedback with their quantitative ratings from Table 2. The inter-rater reliability was assessed using the intraclass correlation coefficient, which yielded a high score of 0.868, indicating strong agreement among raters.
The results shown in Figure D1 highlight that participants’ open-ended sentiments describe the Ignorant agent as a poor teammate. Specifically, the Ignorant agent is represented negatively on teaming phrasing related to Q1 of the questionnaire. The “teaming” construct is frequently mentioned across descriptions of all the agents. On the other hand, the open-ended responses infrequently mentioned terms related to Q6, Q7, and Q8 which were predictive features in the regression model for the choice data (see Figure 8 in the main paper). These results suggests that only a few salient dimensions such as teaming might be reported in the open-ended responses.
Figure D1: Sentiments for each agent based on the open-ended response section in relation to themes from the questionnaire items. Each statement contributed a single count to either positive or negative sentiments. Each open-ended statement was systematically coded as either being positive (green bars) or negative (red bars).
Appendix E Regression Modeling Results
Table E1: Posterior Summaries of Coefficients Predicting Choice in Experiment 1
95% Credible Interval
Model & Coefficient BFinclusion Mean SD Lower Upper
Target Density = 5, Objective Metrics
    Bias $1.000$ $1.386\times 10^{-4}$ $0.131$ 0.262-0.262 $0.246$
    Human Score $1.213$ $0.258$ $0.397$ 0.281-0.281 $1.076$
    AI Score $1.433$ $0.325$ $0.431$ 0.196-0.196 $1.221$
    Score Inequality $7.108$ 0.300-0.300 $0.173$ 0.577-0.577 $0.000$
    AI Steals $8101.536$ 0.745-0.745 $0.190$ 1.105-1.105 0.349-0.349
    Intersections $0.978$ 0.047-0.047 $0.101$ 0.315-0.315 $0.083$
Target Density = 5, Subjective Metrics
    Bias $1.000$ $0.140$ $0.177$ 0.193-0.193 $0.463$
    Team (Q1) $7407.462$ $1.338$ $0.359$ $0.683$ $2.009$
    Competence (Q2) $1.372$ 0.165-0.165 $0.262$ 0.769-0.769 $0.167$
    Understand Bot Intent (Q3) $2.949$ $0.393$ $0.369$ 0.124-0.124 $1.076$
    Understand Human Intent (Q4) $5.006$ $0.500$ $0.369$ 0.003-0.003 $1.159$
    Human Contributed More (Q5) $2.749$ 0.305-0.305 $0.285$ 0.855-0.855 $0.020$
    Easy To Play With (Q6) $0.917$ $0.011$ $0.224$ 0.436-0.436 $0.560$
    Fun To Play With (Q7) $1.087$ $0.123$ $0.320$ 0.344-0.344 $0.922$
    Similar Playing Style (Q8) $5291.962$ $1.264$ $0.337$ $0.642$ $1.901$
Target Density = 15, Objective Metrics
    Bias $1.000$ $0.233$ $0.128$ 0.007-0.007 $0.491$
    Human Score $1.329$ $0.131$ $0.174$ 0.069-0.069 $0.512$
    AI Score $0.768$ $0.058$ $0.143$ 0.172-0.172 $0.448$
    Score Inequality $272.313$ 0.468-0.468 $0.146$ 0.748-0.748 0.199-0.199
    AI Steals $2.454$ 0.201-0.201 $0.186$ 0.580-0.580 $0.000$
    Intersections $0.608$ $0.004$ $0.075$ 0.159-0.159 $0.221$
Target Density = 15, Subjective Metrics
    Bias $1.000$ $1.074$ $0.253$ $0.574$ $1.549$
    Team (Q1) $1.422$ 0.026-0.026 $0.256$ 0.706-0.706 $0.463$
    Competence (Q2) $1.513$ $0.065$ $0.242$ 0.444-0.444 $0.627$
    Understand Bot Intent (Q3) $5.216$ $0.424$ $0.349$ 0.113-0.113 $1.094$
    Understand Human Intent (Q4) $41.769$ $0.817$ $0.384$ $0.000$ $1.466$
    Human Contributed More (Q5) $1.595$ $0.093$ $0.212$ 0.314-0.314 $0.616$
    Easy To Play With (Q6) $290.561$ $1.008$ $0.366$ $0.333$ $1.753$
    Fun To Play With (Q7) $239769.399$ $1.720$ $0.459$ $0.792$ $2.580$
    Similar Playing Style (Q8) $28.285$ $0.689$ $0.334$ $0.000$ $1.262$
Appendix F Additional Results for Experiment 2
Figure F1: Mean proportional score by agent type, AI type and density condition, with standard error bars for Experiment 2. The gray violins demonstrate the distribution of data points.
Figure F2: Mean proportional team score (human + AI) by AI type and density condition with standard error bars for Experiment 2. The overlaid gray violins show how data points are distributed.
Figure F3: Mean questionnaire scores by AI agent type and target density in Experiment 2. Questions were rated on a 7-point Likert-scale. Error bars indicate the standard error from the mean.
Figure F4: Choice preferences across pairs of agents for each target density condition in Experiment 2. Each matrix cell indicates the percentage of participants preferring the row-associated agent A over the column-associated agent B. The side column shows the overall preference percentage for each row agent across all pairings. Asterisks denote choice percentages that significantly deviate from 50%, indicated by a Bayes factor greater than 10. Note that results here are averaged across different presentation orders of the agents (e.g., agent A could have been presented first or second in the experiment).